1.) 4+5-9=
2.) 26-(-9)-6=
3.) 5-5-5=
4.) -4-4+(-15)=

Answer by stanbon(48545)

1.) 4+5-9= 9-9 = 0
------------
2.) 26-(-9)-6= 26+9-6 = 29
------------
3.) 5-5-5= 0-5 = -5
--------------
4.) -4-4+(-15)= -8-15 = -23
===============
Cheers,
Stan H.